OpenStudy (anonymous):

Identify the slope and the y-intercept of the equation: y = -2/3x – 2

hero (hero):

slope = -2/3 y-int = -2

hero (hero):

General form is y = mx+b, where m = slope and b = y-intercept

OpenStudy (anonymous):

ohhh okay, thank you :3

hero (hero):

You can thank me by giving a medal :D

OpenStudy (anonymous):

I do have another question about that if you don't mind helping?